# Cross-Domain Federated Semantic Communication with Global Representation Alignment and Domain-Aware Aggregation

### v1

@@ I Introduction
The number of connected devices is growing at an exponential pace, and this trend is expected to continue at a faster pace due to the increasing demand for seamless connectivity among users [10856890, 10929033]. This rapid proliferation of new devices, coupled with emerging bandwidth-intensive wireless services, will strain the capacity of next-generation wireless systems. Indeed, under a limited bandwidth, next-generation wireless networks must serve a much higher density of users with stricter quality-of-service requirements, such as extremely low latency for online gaming and resource-intensive communication for augmented reality or digital twin applications[11193869, 11192484, qiao2025deepseek, 10930485]. To accommodate these new devices and applications, various communication techniques have been proposed, including the integration of sensing and communication [9724245], mmWave, intelligent networking with the assistance of AI/ML technologies[8755300]. Despite tremendous efforts to improve the communication efficiency of traditional communication, its capacity is limited by Shannon’s theory, thus making it difficult to cope with the emerging transition. Consequently, the research community is shifting toward the concept of semantic communication [10554663], which operates beyond the bit-level paradigm.
@@
Instead of focusing solely on bit reconstruction as done in conventional communication systems, semantic communications aims to successfully interpret the meaning of the message. Semantic communication was introduced by Shannon and Weaver back in the 1950s [shannon1998mathematical], but it did not achieve promising results due to the technology’s limitations. Recently, with the rise of deep learning (DL), semantic communication has gained significant attention due to the feasibility of semantic extraction by DL models. Specifically, researchers have deployed the DL models as the joint source-channel encoder/decoder in the semantic communication system, replacing traditional communication. The DL-based semantic communication has demonstrated outstanding performance in reducing the communication bandwidth, while overcoming the “cliff effect” problem encountered in traditional systems [DJSCCwithfeedback]. The “cliff effect” refers to the failure in transmission when the channel quality decreases below a threshold. On the other hand, the semantic communication performance is robust to channel noise and improves as the noise decreases. The improvements are the result of DL models being trained on an extensive amount of data, which is normally collected and gathered at the centralized server. However, data transmission to a centralized server can lead to privacy concerns due to potential leakage of sensitive information. Therefore, federated learning (FL) has been introduced as a distributed learning framework without the need for centralizing the data [mcmahan2017communication].
@@
In FL, a central server actively and iteratively sends the model to be trained to the devices, i.e., the clients, that possess the data. The clients train the model using their local data and transmit the updated models back to the server, which then aggregates them into a global model to achieve better generalization performance[qiao2025towards]. Several studies have been proposed to enhance the convergence performance of FL with scheduling policies [10041216], device-to-device communication links [9562522], or partial model personalization [pillutla2022federated], whose efforts aim to alleviate the communication resource requirements and enable its deployment in the real world. Existing works have applied the FL framework to train semantic communication models for various tasks, including audio reconstruction [9685654], image reconstruction [10288204, 10531097, wu2025personalized], classification tasks [10937229, 10559618, FLfortaskoriented].
@@
Motivated by this research gap, the main contribution of this paper is the proposal of a novel cross-domain FL framework to train the DL-based joint source-channel encoder/decoder in the semantic communication system. To address the performance degradation of the cross-domain clients in FL[10203389], we first design a tailored mechanism for the server to achieve a generalization performance by aligning the specialized domain knowledge of clients with global knowledge. Secondly, due to the nature of data domains, some domains have a larger number of data samples, while others are more challenging to obtain, resulting in extremely heterogeneous data samples for clients. This unbalanced data across domains can cause the global model to favor the domain with a large number of samples, thus resulting in poor performance for the remaining domains. Therefore, we tailor an aggregation approach to mitigate this degradation in overall performance. Unlike other studies that consider a classification task [wang2025federated], our scenario’s objective is to reconstruct the original image by training the DL-based semantic communication model. The contributions of our study are provided as follows.
@@ II-A Semantic Communication for Image Reconstruction
The need for image transmission has never been reduced and continues to increase significantly throughout history, which motivates extensive research on semantic communication for image reconstruction tasks[DJSCCnofeedback, DJSCCAttention, DJSCCSwin, DJSCCDigital]. The authors in [DJSCCnofeedback] were among the first to recognize the potential of the DL model for end-to-end joint source-channel coding, and they proposed two convolutional neural network models for encoder and decoder. Their proposal achieved higher image quality performance when the channel had low signal-to-noise ratio (SNR) values. However, the paper failed to leverage feedback information for the learning process, which motivated the framework in [DJSCCwithfeedback]. Specifically, they exploited the feedback by analyzing the distortion of the transmitted signal, thereby improving the encoding ability of the DL-based model and achieving higher reconstruction quality. On the other hand, the Attention DL-based JSCC (ADJSCC) was proposed to address the dynamics of the channel state information by exposing the DL models to a wide range of SNR values for parameter optimization [DJSCCAttention]. With the development of the architecture for the DL models, the Swin Transformer [liu2021swin] was proposed to resolve the high computing complexity of the attention mechanism, and thus was adopted in [DJSCCSwin] for securing a more efficient system in both computing and communication resources in the semantic communication system. Approaching semantic communication from a different perspective, the authors in [DJSCCDigital] proposed a novel demodulation method to address the dynamic output of the encoder/decoder, thereby bridging the gaps between semantic communication studies and current digital communication systems. However, the aforementioned studies assume the availability of data on a central server for training and overlook the existence of cross-domain data.
@@ II-C Federated Learning for Semantic Communication
There is an increase in the number of studies on training semantic communication systems with the FL framework, particularly for image transmission [FLSemantic, huh2025feature, FLforGAIassistedSemCom, 11096603, zhou2025fedsem]. Specifically, the authors in [FLSemantic] designed a hybrid model combining convolution neural network (CNN) and Transformer designs for DL-based semantic communication and trained it using FL. To actively update the learning model for the system, the authors divided the clients into two groups: one for full parameter updates and one for aligning feature representation, which stabilizes the optimization process [huh2025feature]. Interestingly, the work [FLforGAIassistedSemCom] adopted generative artificial intelligence to assist the FL training of semantic communication by distilling knowledge to the student model, which ensures a unified update direction for the training. In the work [11096603], the author leveraged cross-modal semantic information to achieve a better user experience, specifically through hybrid coding for the haptic signal. On the other hand, in [zhou2025fedsem], the authors formulated a problem to jointly optimize the energy consumption and model performance of the federated learning framework by controlling the sub-carrier allocation variable, transmission power, and computing frequency. However, none of the above studies considered the domain drift phenomenon when client data differ across domains, which leads to significant degradation in the performance of the learning system.
@@ III-A Semantic Communication Components and Training
In Fig. 2a, we consider the semantic communication model that contains the D-JSCC encoder for the transmitter, a channel model for the wireless environment, and finally a D-JSCC decoder for decoding the signal at the receiver, similar to other semantic studies [DJSCCAttention, DJSCCwithfeedback]. With a given image $I\in$ $\mathbb{R}^{3\times H\times W}$ and a set $\mathcal{S}$ of SNR values, the encoding process of it can be denoted as
@@
where the $\mathrm{Enc}(\cdot)_{\alpha}$ represents the joint source-channel encoder with its corresponding parameters $\alpha$, $X_{I}\in C^{k}$ is the transmitted signal of the image $I$ through the channel. The compression ratio between the transmitted signal and the original image is the ratio between $k$ and $n$, where $n=$ $3\times H\times W$. The signal is exposed not only to noise but also to fading effects caused by obstacles in the wireless environment as follows:
@@ III-B FL for DL-based Semantic Communication System
In (6), the contribution of each local client is proportional to the amount of its data relative to the total data of all clients $D$, which is the core of FedAvg [mcmahan2017communication]. We consider the diversity of clients’ data, reflected across domains such as real-world photos versus cartoon images. The latent space, characteristics, and underlying distributions of images from different domains are distinct, leading to the creation of derived local models and the degradation of the global model.
@@ IV Global Representation and Domain-aware Aggregation for Cross-domain FL
We first define $u$ as the source domain of a client, and the total possible set of source domains is $\mathcal{U}=\{u_{1},u_{2},\ldots,u_{n}\}$. Under the conventional training process of the FL framework, each client trained the received global model with its local data without being aware of the existence of other domains’ data. Therefore, the clients from different domains will update the parameters accordingly to their own data domain regardless of the training process of the other clients, as shown in Fig. 3. This independent updating results in client parameter drift across the network [chen2024fair]. With each client’s final local parameters presented by $\theta_{m}$, and under the domain heterogeneity, the difference between the global and the clients becomes even more significant, which is quantified by:
@@ IV-A Global Representation Feature and Alignment Loss
This generalization loss aligns local data features of a specific domain $S_{m}$ with the global representation across all domains. Compared to a scenario where the local model is trained with a single loss value, the generalization loss provides a soft constraint on parameter updates, inspired by device-level systems constraint in [li2020federated]. To be specific, when the local model’s parameters are trained toward the optimal value for its own domain data, it loses generalization capacity, the generalization loss becomes significantly large, and it will steer the learning process toward a more balanced position. Our loss is combined with the loss of the image reconstruction task to achieve high performance for the general data, regardless of the domain as follows:
@@ IV-B Domain-aware Aggregation for Global Model
Secondly, we identify the dominance problem in the global model aggregation process of the conventional FL framework, where a client with a large amount of data in a specific domain pushes the global model toward its own distribution [mohri2019agnostic]. In FedAvg-like algorithms, the global parameters are updated as a weighted average of local models, where the weights correspond to the data sizes of each client. As a result, clients possessing more abundant data dominate the aggregation, overshadowing the contribution of others. It is worth noting that in real-world scenarios, data from different domains can vary greatly in both availability and complexity. For instance, an artist may require hours or even months to produce a single artwork, while sketch images can be generated within minutes and in large quantities. Consequently, the server may overfit to the easily generated but less semantically rich domain, compromising the generalization of the global model to more meaningful yet underrepresented domains. This imbalance in data samples across domains underscores the necessity of a domain-aware aggregation strategy that accounts semantic alignment and domain diversity rather than solely relying on data volume. To address this domain dominance, we propose a simple yet effective domain-aware mechanism to improve the model’s generalization performance. Specifically, on the server, we first aggregate the models from clients in the same domain to build a specialized model for that domain, utilizing domain-specific information that is general and can be shared with the server. We have a total of four domains: photo, art painting, cartoon, and finally sketch, as shown in Fig. 4. The specialized model for the domain $u$ is presented as follows:
@@
With our proposed aggregation approach, clients with a large amount of data can contribute a greater portion of their trained weights to the specialized models for their corresponding domain, which is reasonable because they are trained on a large amount of data and thus have more advanced models. On the other hand, our approach puts constraints on the weighted contribution of clients from one domain toward a different domain, which prevents the dominance of a high-sample domain in the global aggregation process[wang2024taming]. This ensures the global model is fair to all domains and provides the system’s generalization performance, as indicated in V-D. The sequential detail of the Cross-domain Federated learning framework is given in Algorithm 1. In the first round, we only transmit the initial global model.
@@ IV-C1 Assumptions
We follow three standard assumptions [li2020federated]: 1) Each local objective function is $L_{1}$-Lipschitz smooth and continuous, 2) the stochastic gradient $g_{m,r}=\nabla\mathcal{L}(\theta_{m,r},\xi_{r})$ is unbiased and its variance bounded by $\sigma^{2}$, 3) the expected stochastic gradient is bounded by $V$. On the other hand, we consider the fourth assumption for local representation extraction function, being $L_{2}$-Lipschitz continuous, which can be represented as the following equation[tan2022fedproto]:
@@
$\mathbb{E}[\mathcal{L}_{(r+1)E+1/2}]\leq\;$ $\mathcal{L}_{rE+1/2}-\Big(\eta-\frac{\mathcal{L}_{1}\eta^{2}}{2}\Big)\sum_{e=1/2}^{E-1}\lVert\nabla\mathcal{L}_{rE+e}\rVert^{2}_{2}$
@@
where $rE$ is the timestep before the global feature representation. $rE+1/2$ is the time step right after the global representation and right before its first local step. The theorem indicates the deviation bound of the local objective function after each communication round, and convergence can be assured if a certain expected one-round decrease (with appropriate $\eta$ and $\lambda$).
Corollary 1. The loss function $\mathcal{L}$ of an arbitrary client monotonously decreases in every communication round when
$\eta_{e^{\prime}}<\frac{2\big(\sum_{e=1/2}^{e^{\prime}}\lVert\mathcal{L}_{rE+e}\lVert^{2}_{2}-\lambda L_{2}EV\big)}{L_{1}\big(\sum_{e=1/2}^{e^{\prime}}\lVert\nabla\mathcal{L}_{rE+e}\lVert^{2}_{2}+E\sigma^{2}\big)},$ (16)
where $e^{\prime}$ runs from the first iteration step till last batch of data, and
$\lambda_{e}=\frac{\lVert\nabla\mathcal{L}_{rE+1/2}\lVert^{2}_{2}}{L_{2}EV}.$ (17)
Thus, the loss function converges. We provide the proofs of Theorem 1 and Corollary in [Supplementary].

### v2

@@ I Introduction
The number of connected devices is growing at an exponential pace, and this trend is expected to continue at a faster pace due to the increasing demand for seamless connectivity among users [11, 30]. This rapid proliferation of new devices, coupled with emerging bandwidth-intensive wireless services, will strain the capacity of next-generation wireless systems. Indeed, under a limited bandwidth, next-generation wireless networks must serve a much higher density of users with stricter quality-of-service requirements, such as extremely low latency for online gaming and resource-intensive communication for augmented reality or digital twin applications[45, 22, 29, 27]. To accommodate these new devices and applications, various communication techniques have been proposed, including the integration of sensing and communication [44], mmWave, intelligent networking with the assistance of AI/ML technologies[3]. Despite tremendous efforts to improve the communication efficiency of traditional communication, its capacity is limited by Shannon’s theory, thus making it difficult to cope with the emerging transition. Consequently, the research community is shifting toward the concept of semantic communication [2], which operates beyond the bit-level paradigm.
@@
Instead of focusing solely on bit reconstruction as done in conventional communication systems, semantic communications aims to successfully interpret the meaning of the message. Semantic communication was introduced by Shannon and Weaver back in the 1950s [31], but it did not achieve promising results due to the technology’s limitations. Recently, with the rise of deep learning (DL), semantic communication has gained significant attention due to the feasibility of semantic extraction by DL models. Specifically, researchers have deployed the DL models as the joint source-channel encoder/decoder in the semantic communication system, replacing traditional communication. The DL-based semantic communication has demonstrated outstanding performance in reducing the communication bandwidth, while overcoming the “cliff effect” problem encountered in traditional systems [12]. The “cliff effect” refers to the failure in transmission when the channel quality decreases below a threshold. On the other hand, the semantic communication performance is robust to channel noise and improves as the noise decreases. The improvements are the result of DL models being trained on an extensive amount of data, which is normally collected and gathered at the centralized server. However, data transmission to a centralized server can lead to privacy concerns due to potential leakage of sensitive information. Therefore, federated learning (FL) has been introduced as a distributed learning framework without the need for centralizing the data [19].
@@
In FL, a central server actively and iteratively sends the model to be trained to the devices, i.e., the clients, that possess the data. The clients train the model using their local data and transmit the updated models back to the server, which then aggregates them into a global model to achieve better generalization performance[28]. Several studies have been proposed to enhance the convergence performance of FL with scheduling policies [7], device-to-device communication links [17], or partial model personalization [26], whose efforts aim to alleviate the communication resource requirements and enable its deployment in the real world. Existing works have applied the FL framework to train semantic communication models for various tasks, including audio reconstruction [36], image reconstruction [25, 21, 41], classification tasks [32, 38, 33].
@@
Motivated by this research gap, the main contribution of this paper is the proposal of a novel cross-domain FL framework to train the DL-based joint source-channel encoder/decoder in the semantic communication system. To address the performance degradation of the cross-domain clients in FL[9], we first design a tailored mechanism for the server to achieve a generalization performance by aligning the specialized domain knowledge of clients with global knowledge. Secondly, due to the nature of data domains, some domains have a larger number of data samples, while others are more challenging to obtain, resulting in extremely heterogeneous data samples for clients. This unbalanced data across domains can cause the global model to favor the domain with a large number of samples, thus resulting in poor performance for the remaining domains. Therefore, we tailor an aggregation approach to mitigate this degradation in overall performance. Unlike other studies that consider a classification task [39], our scenario’s objective is to reconstruct the original image by training the DL-based semantic communication model. The contributions of our study are provided as follows.
@@ II-A Semantic Communication for Image Reconstruction
The need for image transmission has never been reduced and continues to increase significantly throughout history, which motivates extensive research on semantic communication for image reconstruction tasks[1, 42, 43, 23]. The authors in [1] were among the first to recognize the potential of the DL model for end-to-end joint source-channel coding, and they proposed two convolutional neural network models for encoder and decoder. Their proposal achieved higher image quality performance when the channel had low signal-to-noise ratio (SNR) values. However, the paper failed to leverage feedback information for the learning process, which motivated the framework in [12]. Specifically, they exploited the feedback by analyzing the distortion of the transmitted signal, thereby improving the encoding ability of the DL-based model and achieving higher reconstruction quality. On the other hand, the Attention DL-based JSCC (ADJSCC) was proposed to address the dynamics of the channel state information by exposing the DL models to a wide range of SNR values for parameter optimization [42]. With the development of the architecture for the DL models, the Swin Transformer [18] was proposed to resolve the high computing complexity of the attention mechanism, and thus was adopted in [43] for securing a more efficient system in both computing and communication resources in the semantic communication system. Approaching semantic communication from a different perspective, the authors in [23] proposed a novel demodulation method to address the dynamic output of the encoder/decoder, thereby bridging the gaps between semantic communication studies and current digital communication systems. However, the aforementioned studies assume the availability of data on a central server for training and overlook the existence of cross-domain data.
@@ II-C Federated Learning for Semantic Communication
There is an increase in the number of studies on training semantic communication systems with the FL framework, particularly for image transmission [15, 10, 24, 8, 46]. Specifically, the authors in [15] designed a hybrid model combining convolution neural network (CNN) and Transformer designs for DL-based semantic communication and trained it using FL. To actively update the learning model for the system, the authors divided the clients into two groups: one for full parameter updates and one for aligning feature representation, which stabilizes the optimization process [10]. Interestingly, the work [24] adopted generative artificial intelligence to assist the FL training of semantic communication by distilling knowledge to the student model, which ensures a unified update direction for the training. In the work [8], the author leveraged cross-modal semantic information to achieve a better user experience, specifically through hybrid coding for the haptic signal. On the other hand, in [46], the authors formulated a problem to jointly optimize the energy consumption and model performance of the federated learning framework by controlling the sub-carrier allocation variable, transmission power, and computing frequency. However, none of the above studies considered the domain drift phenomenon when client data differ across domains, which leads to significant degradation in the performance of the learning system.
@@ III-A Semantic Communication Components and Training
In Fig. 2a, we consider the semantic communication model that contains the D-JSCC encoder for the transmitter, a channel model for the wireless environment, and finally a D-JSCC decoder for decoding the signal at the receiver, similar to other semantic studies [42, 12]. With a given image $I\in$ $\mathbb{R}^{3\times H\times W}$ and a set $\mathcal{S}$ of SNR values, the encoding process is denoted
@@
where the $\mathrm{Enc}(\cdot)_{\alpha}$ represents the joint source-channel encoder with its corresponding parameters $\alpha$, $X_{I}\in C^{k}$ is the transmitted signal of the image $I$ through the channel. The compression ratio between the transmitted signal and the original image is the ratio between $k$ and $n$, where $n=$ $3\times H\times W$. The signal is exposed not only to noise but also to fading effects caused by obstacles in the wireless environment:
@@ III-B FL for DL-based Semantic Communication System
In (6), the contribution of each local client is proportional to the amount of its data relative to the total data of all clients $D$, which is the core of FedAvg [19]. We consider the diversity of clients’ data, reflected across domains such as real-world photos versus cartoon images. The latent space, characteristics, and underlying distributions of images from different domains are distinct, leading to the creation of derived local models and the degradation of the global model.
@@ IV Global Representation and Domain-aware Aggregation for Cross-domain FL
We first define $u$ as the source domain of a client, and the total possible set of source domains is $\mathcal{U}=\{u_{1},u_{2},\ldots,u_{n}\}$. Under the conventional training process of the FL framework, each client trained the received global model with its local data without being aware of the existence of other domains’ data. Therefore, the clients from different domains will update the parameters accordingly to their own data domain regardless of the training process of the other clients, as shown in Fig. 3. This independent updating results in client parameter drift across the network [4]. With each client’s final local parameters presented by $\theta_{m}$, and under the domain heterogeneity, the difference between the global and the clients becomes even more significant, which is quantified by:
@@ IV-A Global Representation Feature and Alignment Loss
This generalization loss aligns local data features of a specific domain $S_{m}$ with the global representation across all domains. Compared to a scenario where the local model is trained with a single loss value, the generalization loss provides a soft constraint on parameter updates, inspired by device-level systems constraint in [16]. To be specific, when the local model’s parameters are trained toward the optimal value for its own domain data, it loses generalization capacity, the generalization loss becomes significantly large, and it will steer the learning process toward a more balanced position. Our loss is combined with the loss of the image reconstruction task to achieve high performance for the general data, regardless of the domain as follows:
@@ IV-B Domain-aware Aggregation for Global Model
Secondly, we identify the dominance problem in the global model aggregation process of the conventional FL framework, where a client with a large amount of data in a specific domain pushes the global model toward its own distribution [20]. In FedAvg-like algorithms, the global parameters are updated as a weighted average of local models, where the weights correspond to the data sizes of each client. As a result, clients possessing more abundant data dominate the aggregation, overshadowing the contribution of others. It is worth noting that in real-world scenarios, data from different domains can vary greatly in both availability and complexity. For instance, an artist may require hours or even months to produce a single artwork, while sketch images can be generated within minutes and in large quantities. Consequently, the server may overfit to the easily generated but less semantically rich domain, compromising the generalization of the global model to more meaningful yet underrepresented domains. This imbalance in data samples across domains underscores the necessity of a domain-aware aggregation strategy that accounts semantic alignment and domain diversity rather than solely relying on data volume. To address this domain dominance, we propose a simple yet effective domain-aware mechanism to improve the model’s generalization performance. Specifically, on the server, we first aggregate the models from clients in the same domain to build a specialized model for that domain, utilizing domain-specific information that is general and can be shared with the server. We have a total of four domains: photo, art painting, cartoon, and finally sketch, as shown in Fig. 4. The specialized model for the domain $u$ is presented as follows:
@@
With our proposed aggregation approach, clients with a large amount of data can contribute a greater portion of their trained weights to the specialized models for their corresponding domain, which is reasonable because they are trained on a large amount of data and thus have more advanced models. On the other hand, our approach puts constraints on the weighted contribution of clients from one domain toward a different domain, which prevents the dominance of a high-sample domain in the global aggregation process[37]. This ensures the global model is fair to all domains and provides the system’s generalization performance, as indicated in V-D. The sequential detail of the Cross-domain Federated learning framework is given in Algorithm 1. In the first round, we only transmit the initial global model.
@@ IV-C1 Assumptions
We follow three standard assumptions [16]: 1) Each local objective function is $L_{1}$-Lipschitz smooth and continuous, 2) the stochastic gradient $g_{m,r}=\nabla\mathcal{L}(\theta_{m,r},\xi_{r})$ is unbiased and its variance bounded by $\sigma^{2}$, 3) the expected stochastic gradient is bounded by $V$. On the other hand, we consider the fourth assumption for local representation extraction function, being $L_{2}$-Lipschitz continuous, which can be represented as the following equation[35]:
@@
$\mathbb{E}[\mathcal{L}_{(r+1)E+\frac{1}{2}}]\leq\;$ $\mathcal{L}_{rE+\frac{1}{2}}-\Big(\eta-\frac{\mathcal{L}_{1}\eta^{2}}{2}\Big)\sum_{e=\frac{1}{2}}^{E-1}\lVert\nabla\mathcal{L}_{rE+e}\rVert^{2}_{2}$
@@
where $rE$ is the timestep before the global feature representation. $rE+\frac{1}{2}$ is the time step right after the global representation and right before its first local step. The theorem indicates the deviation bound of the local objective function after each communication round, and convergence can be assured if a certain expected one-round decrease (with appropriate $\eta$ and $\lambda$). The sketch of proof is as follows:
Lemma 1. With the assumptions 1 and 2 hold, the loss function of a client for communication round $r+1$ to the last local update step can be bounded as:
$\mathbb{E}[\mathcal{L}_{(r+1)E}]\leq\mathcal{L}_{rE+\frac{1}{2}}-(\eta-\frac{L_{1}\eta^{2}}{2})\sum_{e=\frac{1}{2}}^{E-1}\lVert\nabla\mathcal{L}_{rE+e}\lVert^{2}_{2}+\frac{L_{1}E\eta^{2}}{2}\sigma^{2}.$
Proof: With the $\theta_{rE+1}=\theta_{rE+\frac{1}{2}}-\eta g_{rE+\frac{1}{2}}$, the Assumption 1 can be written:
$\mathcal{L}_{rE+1}\leq\mathcal{L}_{rE+\frac{1}{2}}-\eta\langle\nabla\mathcal{L}_{rE+\frac{1}{2}},g_{rE+\frac{1}{2}}\rangle+\frac{L_{1}}{2}\lVert\eta g_{rE+\frac{1}{2}}\lVert^{2}_{2}.$
We take the expectation of both sides on the random variable $\xi_{rE+\frac{1}{2}}$, which we obtain
$\mathbb{E}[\mathcal{L}_{rE+1}]\leq\mathcal{L}_{rE+\frac{1}{2}}-\eta\mathbb{E}[\langle\nabla\mathcal{L}_{rE+\frac{1}{2}},g_{rE+\frac{1}{2}}\rangle]$
$+\frac{L_{1}\eta^{2}}{2}\mathbb{E}[\lVert g_{rE+\frac{1}{2}}\lVert^{2}_{2}],$ (18)
which we then consider the unbiased stochastic gradient from Assumption 2 ($\mathbb{E}_{\xi_{m}\sim D_{m}}=\nabla\mathcal{L}(\theta_{m,r})=\nabla\mathcal{L}_{r}$), $Var(x)=\mathbb{E}[x^{2}]-(\mathbb{E}[x])^{2}$, and finally the variance bounded of Assumption 2 ($\mathbb{E}[\lVert g_{m,r}-\nabla\mathcal{L}(\theta_{m,r})\lVert^{2}_{2}]\leq\sigma^{2}$) , which result in:
$\mathbb{E}[\mathcal{L}_{rE+1}]\leq\mathcal{L}_{rE+\frac{1}{2}}-(\eta-\frac{L_{1}\eta^{2}}{2})\lVert\nabla\mathcal{L}_{rE+\frac{1}{2}}\lVert_{2}^{2}+\frac{L_{1}\eta^{2}}{2}\sigma^{2}.$
Finally, we telescope of E iterations to derive Lemma 1.
Lemma 2. Consider the assumptions 3 and 4 hold, after the global representation aggregation at the server, the loss function for a random client can be bounded as:
$\mathbb{E}[\mathcal{L}_{(r+1)E+\frac{1}{2}}]\leq\mathcal{L}_{(r+1)E}+\lambda L_{2}\eta EV.$ (20)
Proof: Starting with our proposed local loss:
$\mathcal{L}_{(r+1)E+\frac{1}{2}}$ $=\mathcal{L}_{(r+1)E}+\lambda\lVert f({\theta}_{(r+1)E})-G_{r+2}\lVert_{2}$
$-\lambda\lVert f(\theta_{(r+1)E})-G_{r+1}\lVert_{2},$ (21)
we adopt the result of the following math $\lVert a-b\lVert_{2}$-$\lVert a-c\lVert_{2}\leq$ $\lVert b-c\lVert_{2}$ to obtain $\mathcal{L}_{(r+1)E+\frac{1}{2}}\leq\mathcal{L}_{(r+1)E}+\lambda\lVert G_{r+2}-G_{r+1}\lVert_{2}$. After which, we replace $G$ with the definition of global representation, $F$ local representation into the equation:
$\mathcal{L}_{(r+1)E+\frac{1}{2}}$ $\leq\mathcal{L}_{(r+1)E}+\lambda\lVert\sum_{m=1}^{M}\frac{1}{MD_{m}}\sum_{i=1}^{D_{m}}(f_{m}(\theta_{m,(r+1)E},I_{i})$
$-f_{m}(\theta_{m,rE},I_{i})\lVert_{2},$ (22)
Then we apply the math $\lVert\sum a_{i}\lVert_{2}\leq$ $\sum\lVert a_{i}\lVert_{2}$ to obtain the $\lVert(f_{m}(\theta_{m,(r+1)E},I_{i})-f_{m}(\theta_{m,rE},I_{i})\lVert_{2}$, whose values $\leq L_{2}\lVert\theta_{m,(r+1)E}-\theta_{m,rE}\lVert_{2}$ (Assumption 4), thus we obtain:
$\mathcal{L}_{(r+1)E+\frac{1}{2}}\leq\mathcal{L}_{(r+1)E}+\lambda L_{2}\eta\sum_{m=1}^{M}\frac{1}{M}\sum_{e=\frac{1}{2}}^{E-1}\lVert g_{m,rE+e}\lVert_{2},$
Then we take expectation of the random variable $\xi$:
$\mathbb{E}[\mathcal{L}_{(r+1)E+\frac{1}{2}}]$ $\leq\mathcal{L}_{(r+1)E}+\lambda L_{2}\eta\sum_{m}^{M}\frac{1}{M}\sum_{e=\frac{1}{2}}^{E-1}\mathbb{E}[\lVert g_{m,rE+e}\lVert_{2}]$ (24)
$\leq\mathcal{L}_{(r+1)E}+\lambda L_{2}\eta EV,$ (25)
The transition from (24) to (25) follows directly from Assumption 3 and (25) establishes the result stated in Lemma 2.
Finally, we take the expectation of $\theta$ on both sides in Lemmas 1 and 2, and then sum them up to derive Theorem 1.
Corollary 1. The loss function $\mathcal{L}$ of an arbitrary client monotonously decreases in every communication round when $(\eta-\frac{L_{1}\eta^{2}}{2})\sum_{e=1/2}^{E-1}\lVert\nabla\mathcal{L}_{rE+e}\lVert_{2}^{2}+\frac{L_{1}E\eta^{2}}{2}\sigma^{2}+\lambda L_{2}EV<0$. Therefore, the value of learning rate $\eta$ is bounded:
$\eta_{e^{\prime}}<\frac{2\big(\sum_{e=\frac{1}{2}}^{e^{\prime}}\lVert\mathcal{L}_{rE+e}\lVert^{2}_{2}-\lambda L_{2}EV\big)}{L_{1}\big(\sum_{e=\frac{1}{2}}^{e^{\prime}}\lVert\nabla\mathcal{L}_{rE+e}\lVert^{2}_{2}+E\sigma^{2}\big)},$ (26)
where $e^{\prime}$ runs from the first iteration step till the last batch of data, and
$\lambda_{e}=\frac{\lVert\nabla\mathcal{L}_{rE+\frac{1}{2}}\lVert^{2}_{2}}{L_{2}EV}.$ (27)
Thus, the loss function converges.